Calculate the average speed of students of your class we who have participated in 100 metres and 200 metres running race write a report
1
Expert's answer
2019-08-14T13:21:41-0400

To calculate the average speed it is necessary to divide the distance traveled by the time it will take to travel.



"V=\frac{d}{t}"


The formula is given by where:


Where

Distance traveled d

time t


You can calculate the distance traveled for the 100-meter run


But whatdo you need? The time it will say.


For example: the fastest runners do so in approximately 10s or less.


The average speed is:

"v_{100}=\frac{d}{t}=\frac{100m}{10s}=10\frac{m}{s}"


Now for 200 meters the same equation can be applied, only that the distance traveled is d = 200m


considering the time approximately of 20s the average speed is:



"V_{200}=\frac{d}{t}=\frac{200m}{20s}=10\frac{m}{s}"


Although in practice you should use the measured times.


In summary, the average speed is greater while the time the tour is made is shorter, and the average speed says that constant speed that a person must have to make the tour at the same time.


Remember that in a race you are at rest then you accelerate and reach a final speed that you maintain until finishing.


That is why the average speed does not really indicate what happens.